A. Carbonyl fluoride, COF2, is an important intermediate used in the production of fluorine-containing compounds. For instance, it is used to make the refrigerant carbon tetrafluoride, CF4 via the reaction 2COF2(g)⇌CO2(g)+CF4(g), Kc=7.80 If only COF2 is present initially at a concentration of 2.00 M, what concentration of COF2 remains at equilibrium?

B. Consider the reaction

CO(g)+NH3(g)⇌HCONH2(g),    Kc=0.820

If a reaction vessel initially contains only CO and NH3 at concentrations of 1.00 M and 2.00 M, respectively, what will the concentration of HCONH2 be at equilibrium?
Answer #1

1.

the reaction is 2COF2(g) <---->CO2(g)+ CF4(g)

let x= concentration of CO2 at equilibrium, hence at equilibrium [CO2]=[CF4]=x and COF2= 2-2x

hence KC= [CO2] [CF4]/ [COF2]2= 7.8 , x2/(2-2x)2= 7.8

x2/(1-x)2*4= 7.8

x2/(1-x)2= 7.8*4= 31.2

taking square root, x/(1-x)= 5.6

x= 5.6-5.6x, 6.6x= 5.6, x= 5.6/6.6=0.85

at equilibrium, COF2= 2-2*0.85=0.3M

2.

CO(g)+NH3(g)⇌HCONH2(g),    Kc= [HCONH2]/[CO][NH3]= 0.82

let x= drop in concentration of CO to reach equilibrium, at equilibrium [CO]=1-x and [NH3]= 2-x and HCONH2=x

hence Kc= x/(1-x)*(2-x)=0.82, when solved using excel, x= 0.544

at equilibrium [HCONH2]= 0.544